OpenStudy (anonymous):

Find the slope of the line passing through the points (–1, 3) and (4, –7)

OpenStudy (anonymous):

are you familiar with the formula y2-y1/x2-x1?

OpenStudy (anonymous):

yes but i though y2 - y1 means y1 hade to be negative?

OpenStudy (texaschic101):

y1 does not have to be negative.....all it is saying is subtract y1 from y2

OpenStudy (anonymous):

3-(-7) -1-4 10/-5 2/-1

OpenStudy (anonymous):

so the answer would be 2/-1?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (-1,3) x1 = -1 and y1 = 3 (4,-7) x2 = 4 and y2 = -7) now we sub.. slope(m) = (-7 - 3) / (4 - (-1) can you take it from here ?

OpenStudy (anonymous):

i think i can give me a sec

OpenStudy (anonymous):

-10/5 right?

OpenStudy (texaschic101):

reduces to what ?

OpenStudy (texaschic101):

-10/5 reduces to what ?

OpenStudy (anonymous):

i dont know

OpenStudy (texaschic101):

do you know how to reduce fractions ?

OpenStudy (anonymous):

That basically what i got

OpenStudy (texaschic101):

-10/5 reduces to -2/1 or just -2

OpenStudy (texaschic101):

Sammy also got -2

OpenStudy (texaschic101):

do you understand manguitar ?

OpenStudy (anonymous):

You could also try http://www.wolframalpha.com/ that might help

OpenStudy (anonymous):

i get it now thanks! to you all

OpenStudy (texaschic101):

YW :)